1A) Find the frequency of the Balmer series (n=2) spectral line involving n=6.

answer: 7.21*10^14 Hz

2A) What is the binding energy of Neon (Z=10 A=20) assuming it has a mass of 20 u? (you may ignore the mass of electrons)

please show work on paper for both answers
Answer #1

1A) f = R' (1 / n12 - 1/n22)

R' = Rydberg constant expressed in energy units = 3.29e15 Hz

n1 = 2 , n2 = 6

f = 3.25e15 (1/4 - 1/36)

f = 7.21e14 Hz

----------------------------------------------------------------------------------

1B) Ne has a mass of 19.992439 u.

it has 10 protons and 10 neutrons

mass defect = 10 *1.007825 + 10*1.008665 – 19.992439 = 0.172461 u

so, binding energy = 0.172461 * 931.5 = 160.6 MeV